# Scalable multi-qubit intrinsic gates in quantum dot arrays

### v1

@@ III.3 Fidelity estimations
where $d$ is the dimension of the system. A helpful simplification is brought by the fact that $\widetilde{U}_{\mathrm{ideal}}$ is diagonal in the computational basis: we only need to focus on the diagonal elements of $\widetilde{U}$ for calculating the fidelity. Substituting in the diagonal elements in Eq. (8), we can compute the trace product by
@@
is the residue phase accumulated in time from the higher-order energy corrections. Using a combination of inequalities, we can estimate a lower bound for the trace-product, hence the gate fidelity by
$\mathrm{F}\geq 1-\frac{2d}{d+1}\,{\max_{n}\lvert\varphi_{n}\rvert}-\frac{4}{d+1}\sum_{n}{(1-r_{nn})}.$ (45)
We derive this inequality in the appendix. From this lower bound, we find two major sources of coherent errors. One is the residue phase $\varphi_{n}$ accumulated in time. The next is the non-unity component $r_{nn}=\lvert\langle n|\mathopen{}n^{\prime}\rangle\rvert^{2}<1$ of the perturbed eigenstate. Specially, we can apply the non-degenerate perturbation theory to find
@@
$\varphi_{n}\simeq\tau\sum_{m\neq n}\frac{\big{\lvert}\sum_{w}J_{w}\langle n|\mathopen{}\xi_{w}\rangle\langle\xi_{w}|\mathopen{}m\rangle\big{\rvert}^{2}}{\varepsilon_{\mathrm{Z},n}-\varepsilon_{\mathrm{Z},m}}$ (46)
@@
$\sum_{n}{(1-r_{nn})}\simeq\sum_{\stackrel{{\scriptstyle n,m}}{{n\neq m}}}\frac{\big{\lvert}\sum_{w}J_{w}\langle n|\mathopen{}\xi_{w}\rangle\langle\xi_{w}|\mathopen{}m\rangle\big{\rvert}^{2}}{(\varepsilon_{\mathrm{Z},n}-\varepsilon_{\mathrm{Z},m})^{2}}.$ (47)
@@
Although both two terms in Eq. (46) and Eq. (47) seem to be of second order in the exchange energy $J$, the total evolution time contributes inversely $\tau\propto 1/J$. This makes $\varphi_{n}$ first order in $J/\varepsilon_{\mathrm{Z}}$ and can be attributed as the dominant source of error. However, it is possible to suppress this type of error by applying additional corrections to the local phase values predicted by the first-order theory. This technique is already seen in two-qubit gates, where we have slightly different local phase corrections for the two dots for better fidelity.
@@
In general, a set of extra phase corrections produce the gate, $\widetilde{U}^{\prime}=\mathrm{e}^{\mkern 1.0mu\mathrm{i}\mkern 1.0mu\delta\Phi}\widetilde{U}$, with
$\delta\Phi=\delta\phi_{g}+\bigoplus_{j}(0,\delta\phi_{j}),$ (48)
for the compensated free phases $\delta\phi_{g}$ and $\{\delta\phi_{i}\}$ similar as those defined in Eq. (14) and Eq. (15). Comparing the fidelity between $\widetilde{U}^{\prime}$ and $\widetilde{U}_{\mathrm{ideal}}$, we only need to change the vector of accumulated phases $\bm{\varphi}=(\varphi_{1},\varphi_{2},\cdots,\varphi_{2^{N_{Q}}})$ to the compensated values
$\bm{\varphi}\to\bm{\varphi}^{\prime}=\bm{\varphi}-\bm{\delta\Phi}=\bm{\varphi}-K\bm{y},$ (49)
where we introduce the coefficient matrix $K$ is for vector $\bm{y}=(\phi_{g},\phi_{1},\phi_{2},\cdots,\phi_{N_{Q}})$. With the gate fidelity bound by the max norm of the vector $\bm{\varphi}^{\prime}$, the optimal extra phase corrections and accumulated phases can be determined using theory of Moore-Penrose Pseudoinverse by
$\bm{y}=K^{+}\bm{\varphi},\quad\bm{\varphi}^{\prime}=(I-K^{+}K)\bm{\varphi},$ (50)
@@
for $K^{+}\equiv(K^{\mkern-1.0mu\mathsf{T}}K)^{-1}K$ uniquely determined by the system dimension. Combined with optimal values for the tunneling coefficients, the accumulated phases from second order energies correction $\delta E^{(2)}_{n}$ may be removed.
@@
To demonstrate the effect of extra phase corrections in practice. We explicitly consider the example of a MQCP gate on linear array of three-qubit (as “1-C-2”). For simplicity, we assume $s_{j}=\mkern 1.0mu\mathrm{i}\mkern 1.0mu\sin(\theta_{j})$ and $t_{i}=\mathrm{e}^{\mkern 1.0mu\mathrm{i}\mkern 1.0mu\eta_{j}}\sin(\theta_{j})$ for the tunneling coefficients for the two bonds $j=1,2$ connected to the center qubit, with exchange energies $J_{1}$ and $J_{2}$. Applying Eq. (46) and Eq. (50), we find $\bm{\varphi}^{\prime}\simeq\varphi_{\max}\times(1,-1,-1,1,-1,1,1,-1)^{\mkern-1.0mu\mathsf{T}}$, with
$\varphi_{\max}=\tau\frac{J_{1}J_{2}}{8\varepsilon_{\mathrm{Z},2}}\sin\left(2\theta_{1}\right)\sin\left(2\theta_{2}\right)\cos\left(\eta_{1}-\eta_{2}\right).$ (51)
The corresponding local phase compensations $\delta\phi_{1},\delta\phi_{2}$ and $\delta\phi_{C}$ are derived in the appendix. The accumulated phase is much smaller under optimal local phase corrections and can be reduced to higher order in $J/\varepsilon_{\mathrm{Z}}$ if the tunneling coefficients further satisfy conditions such as $\theta_{j}=k\pi/2$ or $\eta_{1}-\eta_{2}=(k+1/2)\pi$ for $k\in\mathbb{Z}$.
III.4 Applications and advantages
@@
The rule of general gate decomposition allows us to conceive other intrinsic multi-qubit gates that can be useful for quantum information processing tasks. Here we construct two simple examples. Hopefully these can inspire more sophisticated discussions in future works.
@@ III.4.1 Three-qubit logical $Z$-gate
The first example is still based on the three-qubit system. But this time we assume all three qubits are all connected to each other, as shown in Fig. 2(b). From the gate decomposition law, the corresponding intrinsic gate can be decomposed as three controlled-phase gates, one for each connecting pair. Assuming that the bonds are homogeneous and that $\pi$-phase flips are applied, the combined gate becomes
@@
As a result, the $|000\rangle$, $|001\rangle$, $|010\rangle$ and $|100\rangle$ states are unaffected while $|011\rangle$, $|101\rangle$, $|110\rangle$ and $|111\rangle$ experience $\pi$-phase (sign) flips. In other words, the resulting phase flips distinguishes the majorly $|0\rangle$ states with the majorly $|1\rangle$ states. We can compare this gate with the simple product of single-qubit gates $Z_{1}Z_{2}Z_{3}$, which flips the sign of the states $|001\rangle$, $|010\rangle$, $|100\rangle$ and $|111\rangle$. Both gates acts the same for $|000\rangle$ and $|111\rangle$, but the later does not respect major voting. Meanwhile, the non-entangling gate $X_{1}X_{2}X_{3}$ correctly interchanges the majorly $|0\rangle$ states with the majorly $|1\rangle$ states and also anticommutes with $G$. Based on this property, this three-qubit system can be considered as a error correction code for bit-flip errors, with the logical $|\bar{0}\rangle$ and logical $|\bar{1}\rangle$ encoded with states that are majorly $|0\rangle$ and majorly $|1\rangle$. The gate $G$ in Eq. (52) becomes the logical $\bar{Z}$ gate whereas $X_{1}X_{2}X_{3}$ becomes the logical $\bar{X}$.
@@ III.4.2 Simultaneous parity checks
The MQCP gates has important implications for parity measurement. The ability to perform simultaneous parity measurement has import applications in quantum error correction codes.
@@ III.4.3 Fast order reversal
Some quantum computing tasks require reversing the order of a set of states. Notably, this step is involved in quantum Fourier transform, which is necessary for the Shor’s algorithm. Here we assume that these states are hosted in a linear array of spin qubits, the quantum circuit for reversing the array typically breaks down into multiple application of swap gates, which is only allowed for near-neighbors due to restriction in connectivity. In general, flipping an $n$-qubit array requires $n(n-1)/2$ near-neighbor swaps, with each swap made up of three CNOT gate (or CZ combined with single-qubit Hadamard gates). Although some swaps can be simultaneously performed, the task still requires $O(n^{2})$ steps to accomplish.
The same array reversal task can be achieved with only $n$ multi-qubit gates with $\pi$-phase flips on all bonds. By the gate decomposition theory, it can be expressed as
$G=(\mathrm{C}_{1}\mathrm{Z}_{2})\cdot(\mathrm{C}_{2}\mathrm{Z}_{3})\cdots(\mathrm{C}_{n-1}\mathrm{Z}_{n}).$ (53)
Consider the following gate sequence that interleaves $n$ such gate with single-qubit Hadamard gates $H=H_{1}\otimes H_{2}\otimes\cdots\otimes H_{n}$,
$R=HGHG\cdots HGH=H(GH)^{n}.$ (54)
Recall that the array reversal map swaps the state $|\bm{a}\rangle$ with $|\overleftarrow{\bm{a}}\rangle$ for all the $n$-digit binary strings $\bm{a}$, e.g., $|0011\rangle\leftrightarrow|1100\rangle$. Using the stabilizer group theory, we can show that $R$ differs from the array reversal map only by single-qubit sign flips depending on the parity of the states, i.e,
$R:\quad|\bm{a}\rangle\leftrightarrow\operatorname{p}(\bm{a})\times|\overleftarrow{\bm{a}}\rangle,$ (55)
where $\operatorname{p}(\bm{a})=\pm 1$ is depends on the number of consecutive $1$ in string $\bm{a}$, with even and odd number of consecutive $1$ producing $p=+1$ and $p=-1$. For example, the even parity strings “$00010$” and “$01110$” contain 0 and 2 occurrence(s) of consecutive $1$, compared with the odd-parity strings “$00110$” and “$01111$” with 1 and 3 occurrence(s). The matrix representation of $R$ is plotted in Fig. 4 for a linear array of 3, 4 and 5 qubits for demonstration. These extra sign flips can further corrected in the software level with virtual $Z$ gates.

### v2

@@ III.3.1 Fidelity bound
where $d$ is the dimension of the system [64]. A helpful simplification is brought by the fact that $\widetilde{U}_{\mathrm{ideal}}$ is diagonal in the computational basis: we only need to focus on the diagonal elements of $\widetilde{U}$ for calculating the fidelity. Substituting in the diagonal elements in Eq. (10), we can compute the trace product by
@@
Using a combination of inequalities, we can estimate a upper bound for the gate infidelity $\mathrm{InF}=1-\mathrm{F}$ by
$\mathrm{InF}\leq\frac{4}{d+1}\sum_{n}{(1-r_{nn})}+\frac{1}{d+1}\sum_{n}{\varphi^{2}_{n}}.$ (47)
We derive this inequality in the Appendix B. From this bound, we identify two contributing factors of coherent errors. One is the decreased overlap between the perturbed and unperturbed eigenstates $r_{nn}\lesssim 1$. The other is the residue phase $\varphi_{n}$ accumulated in time from the higher-order energy corrections. The magnitudes of these terms can be estimated by carrying out second-order perturbation. The non-unit wave function overlap
$1-r_{nn}\simeq\sum_{m\neq n}\frac{\big{\lvert}\langle n|H_{\mathrm{ex}}|m\rangle\big{\rvert}^{2}}{(E_{n}-E_{m})^{2}}$ (48)
contributes to the infidelity of second-order in the perturbative strength $J/\varepsilon^{\mathrm{Z}}$. While the accumulated phase
@@
$\varphi_{n}$ $\simeq\tau\sum_{m\neq n}\frac{\big{\lvert}\langle n|H_{\mathrm{ex}}|m\rangle\big{\rvert}^{2}}{E_{n}-E_{m}}$ (49)
appears quadratic in $H_{\mathrm{ex}}$, the total evolution time $\tau$ is reversely proportional to $J$ for typical quantum gates. Therefor the second term in Eq. (47) is also of second order in $J/\varepsilon^{\mathrm{Z}}$.
The estimations in Eq. (48) and Eq. (49) make sense only when the eigenenergies of $H_{\mathrm{qu}}$ are non-degenerate. Indeed, as $\widetilde{U}_{\mathrm{ideal}}$ is defined from the non-degenerate first-order perturbation, we expect the approximation $\widetilde{U}(\tau)\approx\widetilde{U}_{\mathrm{ideal}}$ to fail if there is energy degeneracy. This is reflected in the diverging infidelity upper bound for the right-hand-side of Eq. (47). Although the degeneracy problem can be circumvent by imposing the additional condition $\lvert E_{m}-E_{n}\rvert>\lVert H_{\mathrm{ex}}\rVert$ for all $m,n$, such requirement is impractical for system with a large number of quantum dots, in which the eigenenergies $\{E_{n}\}$ become densely packed. Therefore, a careful analysis accounting for accidental degeneracies is necessary.
A closer inspection of Eq. (48) and Eq. (49) reveals that it is not necessary to require $E_{n}\neq E_{m}$ for all $m\neq n$. In most cases, the numerator in the sum vanishes and does not contribute to gate error. The matrix element $\langle n|H_{\mathrm{ex}}|m\rangle$ is non-zero only when the Hamming distance between $m$ and $n$ is less or equal to $2$. In other words, for fixed $|n\rangle=|n_{1},n_{2}\cdots,n_{N_{Q}}\rangle,$ with spin component $n_{j}\in\{\uparrow,\downarrow\}$, we only need to consider the $|m\rangle$ states whose spin components differ from that of $|n\rangle$ in at most two locations. This is evident from the fact that $H_{\mathrm{ex}}$ is the Kronecker sum of four-dimensional matrices. Furthermore, for $\langle n|H_{\mathrm{ex}}|m\rangle\neq 0$, the two dots corresponding to different spin components in $|n\rangle$ and $|m\rangle$ should be connected (with nonzero $J$ coupling). As a result, we can reduce the infidelity bound into a summation of error contributions from each bond
@@
$\sum_{n}(1-r_{nn})\simeq\sum_{w}J^{2}_{w}\,\sum_{\mathclap{n_{w},m_{w}}}\frac{\lvert\langle n_{w}|\mathopen{}\xi_{w}\rangle\langle\xi_{w}|\mathopen{}m_{w}\rangle\rvert^{2}}{(E_{n_{w}}-E_{m_{w}})^{2}}$ (50)
$\sum_{n}\varphi_{n}^{2}\lesssim 2\sum_{w}\tau^{2}J_{w}^{4}\,\sum_{\mathclap{n_{w},m_{w}}}\frac{\lvert\langle n_{w}|\mathopen{}\xi_{w}\rangle\langle\xi_{w}|\mathopen{}m_{w}\rangle\rvert^{4}}{(E_{n_{w}}-E_{m_{w}})^{2}},$ (51)
where $|n_{w}\rangle$ and $|m_{w}\rangle$ is the projection of $|n\rangle$ and $|m\rangle$ in the four-dimensional subspace of the bond $w$. For $w=\langle j,k\rangle$, the possible denominators in Eq. (50) and Eq. (51) are ±$\pm\frac{1}{2}(\varepsilon^{\mathrm{Z}}_{j}+\varepsilon^{\mathrm{Z}}_{k})$ and $\frac{1}{2}(\varepsilon^{\mathrm{Z}}_{j}\pm\varepsilon^{\mathrm{Z}}_{k})$. Therefore it is sufficient to require the Zeeman energies between neighboring dots to be non-degenerate, $\lvert\varepsilon^{\mathrm{Z}}_{j}-\varepsilon^{\mathrm{Z}}_{k}\rvert>J_{jk}$—a much weaker condition than requiring all $E_{n}$ to be non-degenerate. Physically, we can require the Zeeman energy splittings to differ across neighboring quantum dots by having significant $g$-factor variations or inhomogeneous magnetic field.
III.3.2 Reducing coherent error
@@
It is possible to reduce the coherent error by applying additional local phase corrections to the values in Eq. (43) predicted by the first-order theory. This technique is already seen in two-qubit gates, where slightly different local phase corrections are applied for the two dots for improved fidelity [27].
@@
We extend such idea to an intrinsic multi-qubit gate. Similar as the free phase map defined in Eq. (16), applying an extra set of local phase corrections to the multi-qubit gate $\widetilde{U}$ produces $\widetilde{U}^{\prime}=\mathrm{e}^{\mkern 1.0mu\mathrm{i}\mkern 1.0mu\delta\Phi}\widetilde{U}$, with
${\delta\bm{\Phi}}=\bigoplus_{j}(0,\delta\phi_{j})\equiv K{\delta\bm{\phi}},$ (52)
for additional phase shift $\delta\phi_{j}$ on dot $j$. For convenience, we introduce the vector ${\delta\bm{\phi}}=(\phi_{1},\phi_{2},\cdots,\phi_{N_{Q}})^{\mkern-1.0mu\mathsf{T}}$ and the coefficient matrix $K$, which is only determined by $N_{Q}$ and can be efficiently calculated. To compute the improved fidelity between $\widetilde{U}^{\prime}$ and $\widetilde{U}_{\mathrm{ideal}}$, we only need to replace the vector of accumulated phases ${\bm{\varphi}}$ with the shifted values
${\bm{\varphi}}\to{\bm{\varphi}^{\prime}}={\bm{\varphi}}-{\bm{\delta\Phi}}={\bm{\varphi}}-K{\delta\bm{\phi}},$ (53)
where the components of $\bm{\varphi}$ is given in Eq. (46). According to Eq. (47), the gate infidelity is contributed by the norm of the vector $\bm{\varphi}$. Therefore, our goal is to find an optimal vector of extra phases ${\delta\bm{\phi}}$ such that the norm of $\bm{\varphi}^{\prime}$ is minimized. Now if $K$ is invertible, it is easy to see that such optimal ${\delta\bm{\phi}}=K^{-1}\bm{\varphi}$. But $K$ cannot be inverted for $n_{Q}\geq 2$ as it is a $2^{N_{Q}}\times N_{Q}$ dimensional. In such case, the optimal extra phase vector and correct phase vector is defined by
${\delta\bm{\phi}}=K^{+}\bm{\varphi},\quad\bm{\varphi}^{\prime}=(I-K^{+}K)\bm{\varphi},$ (54)
@@
for the Moore-Penrose pseudoinverse $K^{+}\equiv(K^{\mkern-1.0mu\mathsf{T}}K)^{-1}K$ of $K$ [65]. Combined with optimal values for the tunneling coefficients, the accumulated phases from second order energies correction $\delta E^{(2)}_{n}$ can be effectively suppressed.
@@
To demonstrate the effect of extra phase corrections in practice. We explicitly consider the example of a MQCP gate on linear array of three-qubit (as “1-C-2”). Without loss of generality, we set $\tilde{s}_{j}=\mkern 1.0mu\mathrm{i}\mkern 1.0mu\sin(\theta_{j})$ and $\tilde{t}_{i}=\mathrm{e}^{\mkern 1.0mu\mathrm{i}\mkern 1.0mu\eta_{j}}\cos(\theta_{j})$ for the tunneling coefficients for the two bonds $j=1,2$ connected to the center qubit, with exchange energies $J_{1}$ and $J_{2}$. Applying applying the optimal local phase correction Eq. (54) to the $\bm{\varphi}$ vector specified by Eq. (49), we find $\bm{\varphi}^{\prime}\simeq\varphi_{\max}\times(1,-1,-1,1,-1,1,1,-1)^{\mkern-1.0mu\mathsf{T}}$, with
$\varphi_{\max}=\tau\frac{J_{1}J_{2}}{8\varepsilon^{\mathrm{Z}}_{2}}\sin\left(2\theta_{1}\right)\sin\left(2\theta_{2}\right)\cos\left(\eta_{1}-\eta_{2}\right).$ (55)
The norm of $\bm{\varphi}^{\prime}$ is much smaller than that of the uncorrected $\bm{\varphi}$. In fact, from Eq. (55) we see that the second order error contribution can be fully reduced if the tunneling coefficients further satisfy conditions such as $\theta_{j}=k\pi/2$ or $\eta_{1}-\eta_{2}=(k+1/2)\pi$ for $k\in\mathbb{Z}$.
III.4 Applications of multi-qubit gates
@@
The rule of general gate decomposition allows us to conceive many intrinsic multi-qubit gates that can be useful for quantum information processing tasks. Here we consider three simple examples of their advantageous applications. Hopefully these discussion can inspire more sophisticated multi-qubit gates in the future.
@@ III.4.1 Three-qubit logical $Z$-gate
Our first example is based on the three-qubit system. But instead of a linear geometry, this time we consider the case where all three qubits are all connected to each other, as shown in Fig. 2(b). From the gate decomposition law, the corresponding intrinsic gate can be decomposed as three controlled-phase gates, one for each connecting pair. Assuming that the bonds are homogeneous and that $\pi$-phase flips are applied, the combined gate becomes
@@
As a result, the $|000\rangle$, $|001\rangle$, $|010\rangle$ and $|100\rangle$ states are unaffected while $|011\rangle$, $|101\rangle$, $|110\rangle$ and $|111\rangle$ experience $\pi$-phase (sign) flips. In other words, the phase flips for $G$ distinguishes the majorly $|0\rangle$ states with the majorly $|1\rangle$ states. Meanwhile, the non-entangling gate $X_{1}X_{2}X_{3}$ correctly interchanges the majorly $|0\rangle$ states with the majorly $|1\rangle$ states and also anticommutes with $G$. Based on this property, this three-qubit system can be considered as a error correction code for bit-flip errors, with the logical $|\bar{0}\rangle$ and logical $|\bar{1}\rangle$ encoded with states that are majorly $|0\rangle$ and majorly $|1\rangle$. The gate $G$ in Eq. (56) becomes the logical $\bar{Z}$ gate whereas $X_{1}X_{2}X_{3}$ becomes the logical $\bar{X}$. In comparison to $G$, we consider the simple product gate of $Z_{1}Z_{2}Z_{3}$, which flips the sign of the states $|001\rangle$, $|010\rangle$, $|100\rangle$ and $|111\rangle$. Both gates anticommute with $X_{1}X_{2}X_{3}$ and act the same on $|000\rangle$ and $|111\rangle$. But the $Z_{1}Z_{2}Z_{3}$ gate does not respect major voting like $G$, therefore it cannot serve as the logical $\bar{Z}$ for our code.
@@ III.4.2 Simultaneous parity checks
The ability to perform parity measurements on different qubits is a fundamental requirement for many quantum error correction codes [66]. It turns out that the MQCP gates are particularly suitable for such parity measurement tasks. We demonstrate some explicit examples of MQCP as parity checkers in Fig. 3.
@@ III.4.3 Fast order reversal
Some quantum computing tasks require reversing the order of a set of states—notably, this step is involved in quantum Fourier transform [66]. Assuming that a multi-qubit state is stored in a linear array of $N$ qubits, the task of order reversal is to “flip” the array such that the basis state $|\bm{n}\rangle$ is swapped with its reversed state $|\overleftarrow{\bm{n}}\rangle$,
$|\bm{n}\rangle=|n_{1},n_{2},\cdots,n_{N}\rangle\leftrightarrow|n_{N},\cdots,n_{2},n_{1}\rangle\equiv|\overleftarrow{\bm{n}}\rangle,$ (57)
e.g., $|0011\rangle\leftrightarrow|1100\rangle$. The map should be linear such that any quantum entanglement is preserved.
The quantum circuit for reversing the array typically breaks down into multiple application of swap gates, which is only allowed for near-neighbors due to restriction in connectivity. In general, flipping an $N$-qubit array requires $N(N-1)/2$ near-neighbor swaps, with each swap made up of three CNOT gate (or CZ combined with single-qubit Hadamard gates). Although some swaps can be simultaneously performed, the task still requires $O(N^{2})$ steps to accomplish. However, such array reversal task can be achieved with only $N$ applications of the intrinsic multi-qubit gate $L$ with $\pi$-phase flips on all bonds. Such gate can be decomposed as the product
$L=(\mathrm{C}_{1}\mathrm{Z}_{2})\cdot(\mathrm{C}_{2}\mathrm{Z}_{3})\cdots(\mathrm{C}_{N-1}\mathrm{Z}_{N}).$ (58)
Consider the following gate sequence that interleaves the $L$ gate with single-qubit Hadamard gates $H=H_{1}\otimes H_{2}\otimes\cdots\otimes H_{N}$,
$R=HLHL\cdots HLGH=H(LH)^{N}.$ (59)
Using the stabilizer group formalism [67], we can show that $R$ differs from the order reversal map only by single-qubit sign flips depending on the parity of the states,
$R:\quad|\bm{n}\rangle\leftrightarrow\operatorname{p}(\bm{n})\times|\overleftarrow{\bm{n}}\rangle,$ (60)
where the parity function $\operatorname{p}(\bm{n})=\pm 1$ depends on the number of consecutive $1$ in string $\bm{n}$, with even and odd number of consecutive $1$ producing $p=+1$ and $p=-1$. For example, the even parity strings “$00010$” and “$01110$” contain 0 and 2 occurrence(s) of consecutive $1$, compared with the odd-parity strings “$00110$” and “$01111$” with 1 and 3 occurrence(s). The matrix representation of $R$ is plotted in Fig. 4 for the examples of a linear array of 3, 4 and 5 qubits for demonstration. These extra sign flips can further corrected in the software level following the above specified rule. As a result, application of the multi-qubit gate accelerate the order reversal task to $O(N)$ time steps, compared with the equivalent circuit of two-qubit gates taking $O(N^{2})$ steps.